Solve for x in the equation
3sin^2x = cos^2x; 0 3 sin squared x = cos squared x where x is greater than or = 0 and less than or = 2pi
Thank you so much for your help!

Answer by stanbon(48516)

Solve for x in the equation
3sin^2x = cos^2x; 0 3 sin squared x = cos squared x where x is greater than or = 0 and less than or = 2pi
------------------------------------
Solve for x in the equation
3sin^2x = cos^2x
---
Divide both sides by 3cos^2x to get:
tan^2(x) = 1/3
tan(x) = 1/sqrt(3)
---
x = pi/6 or (7/6)pi
=============================
Cheers,
Stan H.